# Reputation in the Shadow of Exit

### v5

@@ Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in the unique Markov equilibrium. Along the way, I highlight the role of Markov strategies in pinning down the value of reputation formation. I apply my results to give qualified commitment foundations for the infinite chain-store game. I also analyze repeated global games with exit, and obtain new predictions about regime survival.
@@ 1 Introduction
Reputational incentives have long been recognized as a powerful way to refine predictions in infinitely repeated games. By endowing the long-run player with the ability to imitate a commitment type who always takes some action $a$, a rational long-run player can build a reputation for playing $a$ and thus secure the payoff they would obtain as-if they were known to be committed to that action. Successful reputation formation relies on two important features of the repeated game: First, that the game itself is repeated sufficiently often, so that the long-run player can undertake the action enough times to build their reputation, and second, that they are extremely patient, so that the “reputation-building” phase of the game contributes little to their average payoff.
@@
Yet there are a variety of settings where the first assumption may be unrealistic. Consider, for example, the chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982) (hereafter, simply “chain store game” or “KWMR”). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position in an a-priori fixed number of markets, even if they unsuccessfully deter entry. Yet in many incumbent-entrant games, successful entry by a challenger can displace the incumbent, changing the structure of the game in future periods, and either forcing the incumbent to exit or the game to prematurely end [^2] For example, the monopolist loses their market power in the KWMR chain store game if a challenger successfully enters and is accommodated.. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in models of repeated regime change (c.f. Huang (2017)), a successful currency attack could end the current regime, i.e. lead to the toppling of the central bank. Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
@@
In such games, the assumption the stage game will be infinitely repeated can be unreasonable, as player actions may lead to premature permanent exit. Consequently, the expected number of periods play persists may be much shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” (c.f. Fudenberg and Levine (1989)) blocking both the long-run player’s ability to build a reputation and for the early stages of the game to be unimportant for their average payoff. Indeed, without strong assumptions on the exit technology itself, little can be said about reputation formation in the shadow of exit, restoring the indeterminancy of the folk theorem present in the complete information analogue (see Proposition 3). Despite the potential appropriateness of directly modeling exit in several of the leading examples studied by the reputation literature, little has been said about the effect exit has on incentives for reputation formation.
@@
This paper studies reputation formation in repeated games where player actions endogenously determine if the game permanently ends. Formally, I consider a model where many short-run players (choosing $a$) interact with a long-run player (choosing $c$) who can be either a rational type or a single committed type who plays a Stackelberg action. In addition to payoffs, the action profile $(a,c)$ also jointly determines a survival probability the game permanently ends. The rational type wishes to maximize their expected payoff by potentially imitating the committed type. Crucially, fear of exit can shut down traditional incentives for reputation formation, as the effective discount rate given short-run player actions can be bounded away from $1$ even when the long-run player is very patient. Thus there is no guarantee that investing in the Stackelberg action today will be rewarded with short-run Stackelberg best replies with high frequency in the future, rendering standard arguments from the reputation literature inapplicable.
@@
    Incentive compatibility: the “effective” discount fact is high enough that playing the Stackelberg action can be enforced in some equilibrium of the repeated game.
@@
Theorem 1 leads to novel comparisons of Markov equilibria. In the complete information game, Markov equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 1 and 2). However, when the regime faces reputational incentives, the unique Markov equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type. [^3] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021) Thus, the Markov refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to strictly outperform their best Markov payoff in the complete information game when they have incentives for reputation formation.
I illustrate my results by returning to the chain store game. I show that when accommodating the entrant leads to some probability the game permanently ends, one of two things can happen depending on whether the credible beliefs restriction is adopted. First, in Markov equilibrium, there exists a unique equilibrium profile which secures the long-run player their Stackelberg payoff. However, in sequential equilibrium, the long-run player’s payoff may secure only their minimax payoff, exactly because of pathological equilibria where short-run players vary play in ways that are payoff-irrelevant. That the long-run player cannot secure a high payoff in sequential equilibrium contrasts sharply with the results of the KWMR chain store game absent exit. I interpret these results as giving qualified commitment foundations for the chain store game with exit.
@@ 1.1 Related Literature
This paper contributes to the extensive literature on when reputational incentives can select payoffs, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the canonical framework of Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide class of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature studying variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, Pei shows that endogenous records accord differing roles to the exit probability and patience of players, and consequently there is rich asymptotic behavior depending on which force in the discount rate is taken to $1$. In contrast, I endogenize the exit probability instead of records, so in particular allow it to always be bounded away from its patient limit. This leads to a different exercise than Pei’s models, and thus different predictions despite a similar decomposition of the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
@@
Using Markov refinements to make predictions in reputational repeated games is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markov refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives, while Kostadinov and Roldán (2025) focus on Markov equilibrium to characterize the dynamics of optimal monetary policy with reputational concerns. Giannitsarou and Toxvaerd (2003) use Markov refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for Markov criteria, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
Finally, the regime change application builds on the rich literature on equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). The model’s key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The first change relates to a large literature in global games where some “large” player can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition B.1 makes an independent contribution in this direction by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique Markov equilibrium.
@@ 2.1 The Stage Game
There are $N+1$ players, $i\in[N]\cup\{L\}$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact set $a_{i}\in A_{i}$; let $a=\{a_{i}\}\in\prod_{i}A_{i}=A$ denote a profile of short-run player strategies. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0.555In the global games application, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 3: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\overline{f}:X\to\mathbb{R}$ such that $\overline{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis. Short-run player actions generate finite signal $Y$, drawn independently according to a distribution $\rho(\cdot|a,c)$. [^6] Short-run player monitoring is irrelevant to the analysis and included only for generality. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, represented by a function $f:A\times C\to[0,1]$. This can be interpreted as the survival probability of a regime given a short-run action profile $a$ and long-run action $c$. The complementarity probability 1f(a,c)1-f(a,c) is the exit probability.
The long-run player’s expected stage payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff they attain conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$. Suppose all payoffs are bounded and continuous in all arguments. Moreover, assume that $f(a,c)<1-\xi$ for some $\xi>0$ always. This can be interpreted as saying there is always some (potentially arbitrarily small) probability exit occurs regardless of what actions are taken. [^7] This ensures the undiscounted payoff streams are well-defined, and in particular converge.
For any long-run player action $c$, denote by $\mathcal{N}(c)$ the Nash correspondence at $c$; that is,
$\mathcal{N}(c)=\left\{a:a_{i}\in\operatorname*{\arg\!\max}_{\overline{a}_{i}\in A_{i}}u_{i}(\overline{a}_{i},a_{-i},c)\right\}.$
@@
$a^{*}(c)=\operatorname*{\arg\!\min}_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
as the set of actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no ambiguity, I also abuse notation and use $a^{*}(c)$ to mean a selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption a (myopic) long-run player who need not worry about exit has a dominant action, labeled at $c=0$, and that they benefit from commitment.
@@
For all $a$, $0=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$. Moreover,
$0\not\in\operatorname*{\arg\!\max}_{c\in C}\min_{a\in A}\left\{f(a,c)v(a,c)\right\}$
The first part of the assumption implies the set of payoffs attainable in the one-shot game can be represented as $\{f(a,0)v(a,0)\}_{a\in\mathcal{N}(0)}$, i.e. the profiles where the long-run player chooses 0 and short-run players best reply. This simplifies the exposition by simplifying the “worst” punishments from deviating from a prescribed equilibrium (though at the cost of substantially more notational overhead, is not necessary). Towards this end, let $\overline{V}_{0}=\max_{a\in\mathcal{N}(0)}\frac{f(a,0)v(a,0)}{1-f(a,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}$ be their smallest such payoff. The second part of the assumption guarantees guarantees the long-run player strictly benefits from the ability to build a reputation when compared to their worst stage-game Nash payoff.

### v6

@@ Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in all Markov equilibrium. Along the way, I highlight the role of Markov strategies in pinning down the value of reputation formation. I apply my results to give conditional commitment foundations for the infinite chain-store game. I also analyze repeated global games with exit, and obtain new predictions about regime survival.
@@ 1 Introduction
Reputational incentives have long been recognized as a powerful way to refine predictions in infinitely repeated games. By endowing the long-run player with the ability to imitate a commitment type who always takes some action $a$, a rational long-run player can build a reputation for playing $a$, and thus secure the payoff they would obtain as-if they were known to be committed to that action. Successful reputation formation relies on two important features of the repeated game: First, that the game itself is repeated sufficiently often, so that the long-run player can undertake the commitment action enough times to build their reputation, and second, that the long-run player is patient relative to the speed of reputation formation, so that the “reputation-building” phase of the game contributes little to the long-run player’s average payoff.
@@
Yet there are a variety of settings where these assumptions may fail to hold. Consider, for example, the chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982) (hereafter, simply “chain store game” or “KWMR”). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position in an a-priori fixed number of markets, even if they unsuccessfully deter entry. Yet in many incumbent-entrant games, including the chain store game, successful entry by a challenger can displace the incumbent, changing the structure of the game in future periods, and either forcing the incumbent to exit or the game to prematurely end [^2] For example, the monopolist loses their market power in the KWMR chain store game if a challenger successfully enters and is accommodated.. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in models of repeated regime change (c.f. Huang (2017)), a successful currency attack could end the current regime, i.e. lead to the toppling of the central bank. Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
@@
In such games, the assumption that the stage game will be infinitely repeated can be unreasonable, as player actions may lead to premature permanent exit. Consequently, the expected number of periods play persists may be much shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” (c.f. Fudenberg and Levine (1989)) blocking both the long-run player’s ability to build a reputation and for the early stages of the game to be unimportant for their average payoff. Indeed, without strong assumptions on the exit technology itself, little can be said about reputation formation in the shadow of exit, restoring the indeterminancy of the folk theorem present in the complete information analogue (see Proposition 3). Despite the potential appropriateness of directly modeling exit in several of the leading examples studied by the reputation literature, little has been said about the effect exit has on incentives for reputation formation.
@@
This paper studies reputation formation in repeated games where player actions endogenously determine if the game permanently ends. Formally, I consider a model where many short-run players (choosing $a$) interact with a long-run player (choosing $c$) who can be either a rational type or a single committed type who plays a Stackelberg action. In addition to payoffs, the action profile $(a,c)$ also jointly determines a survival probability that the game continues; with complementary probability, the game permanently ends. The rational type wishes to maximize their expected payoff by potentially imitating the committed type. Crucially, fear of exit can shut down traditional incentives for reputation formation, as the effective discount rate given short-run player actions can be bounded away from $1$ even when the long-run player is very patient. Thus there is no guarantee that investing in the Stackelberg action today will be rewarded with short-run Stackelberg best replies with high frequency in the future, rendering standard arguments from the reputation literature inapplicable.
@@
    Incentive compatibility: the “effective” discount factor is high enough that playing the Stackelberg action can be enforced in some equilibrium of the repeated game.
@@
Theorem 1 leads to novel comparisons of Markov equilibria. In the complete information game, Markov equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 1 and 2). However, when the regime faces reputational incentives, the unique Markov equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type [^3] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021). Thus, the Markov refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to strictly outperform their best Markov payoff in the complete information game when they have incentives for reputation formation.
I illustrate my results by returning to the chain store game. I show that when accommodating the entrant leads to some probability the game permanently ends, one of two things can happen depending on whether the credible beliefs restriction is adopted. First, in Markov equilibrium, there exists a unique equilibrium profile which secures the long-run player their Stackelberg payoff. However, in sequential equilibrium, the long-run player’s payoff may secure only their minimax payoff, exactly because of pathological equilibria where short-run players vary play in ways that are payoff-irrelevant. That the long-run player cannot secure a high payoff in sequential equilibrium contrasts sharply with the results of the chain store game absent exit. I interpret these results as giving qualified commitment foundations for the chain store game with exit.
@@ 1.1 Related Literature
This paper contributes to the extensive literature on reputation formation in repeated games, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the canonical framework of Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide class of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature studying variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, Pei shows that endogenous records accord differing roles to the exit probability and patience of players, and consequently there is rich asymptotic behavior depending on whether discounting or exit vanishes first. In contrast, I endogenize the exit probability instead of records, so in particular allow it to always be bounded away from its patient limit. This leads to a different exercise than Pei’s models, and thus different predictions despite a similar decomposition of the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
This paper also relates to a smaller literature on strategy-contingent discounting in repeated games without reputation effects. Koppel (2004) studies the prisoner’s dilemma in a complete information game with endogenous discounting, while Neilson and Winter (1996) give a general folk theorem when the discount rate is endogenously chosen by one player. When player payoffs are not time-separable, including as a special case the setting where the discount rate is endogenous, Kochov and Song (2023) and Kochov and Song (2025) present a general folk theorem. I contribute to this literature by (1) studying a specific form of time endogenous discounting—permanent exit in the game, and (2) analyzing the perturbed game with reputation effects.
@@
Using Markov refinements to make predictions in repeated games with reputation effects is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markov refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives, while Kostadinov and Roldán (2025) focus on Markov equilibrium to characterize the dynamics of optimal monetary policy with reputational concerns. Giannitsarou and Toxvaerd (2003) use Markov refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for the Markov refinement, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
Finally, the regime change application builds on the rich literature of equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). The model’s key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The first change relates to a large literature in global games where some “large” player can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition B.1 makes an independent contribution in this direction by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique Markov equilibrium.
@@ 2.1 The Stage Game
There are $N+1$ players in each period, $i\in\{L\}\cup[N]$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact topological space $a_{i}\in A_{i}$; let $a\in\prod_{i}A_{i}=A$ denote a profile of short-run player (pure) actions. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0555In the global games application, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 3: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\overline{f}:X\to\mathbb{R}$ such that $\overline{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis.. The entire action profile also generates a finite signal $Y$, drawn according to a distribution $\rho(\cdot|a,c)$ [^6] Short-run player monitoring is irrelevant to the analysis and included only for generality.. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, represented by a function $f:A\times C\to[0,1]$. This can be interpreted as the probability that the game continues. The complementarity probability 1f(a,c)1-f(a,c) is the exit probability.
The long-run player’s expected stage game payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff they attain conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$ [^7] Modifying short-run player payoffs to be $f(a,c)u_{i}(a,c)$ will not affect any of the analysis, though it will implicitly change which actions are best replies and thus implicitly affect enforceability constraints.. Suppose all payoffs are bounded and continuous in all arguments. Moreover, assume that $f(a,c)<1-\xi$ for some $\xi>0$ always. This can be interpreted as saying there is always some (arbitrarily small) probability exit occurs regardless of what actions are taken [^8] This ensures the undiscounted payoff streams are well-defined, and in particular converge..
For any (potentially mixed) long-run player action $\gamma\in\Delta(C)$, denote by $\mathcal{N}(\gamma)$ the Nash correspondence at $\gamma$ over all strategies; that is,
$\mathcal{N}(\gamma)=\left\{\alpha\in\prod_{i}\Delta(A_{i}):\text{supp}(\alpha_{i})\subset\operatorname*{\arg\!\max}_{\overline{a}_{i}\in A_{i}}\mathbb{E}_{a_{-i}\sim\alpha_{-i}}\mathbb{E}_{c\sim\gamma}[u_{i}(\overline{a}_{i},a_{-i},c)]\right\}.$
@@
$a^{*}(c)=\operatorname*{\arg\!\min}_{\alpha\in\mathcal{N}(c)}\left\{\mathbb{E}_{a\sim\alpha}[f(a,c)v(a,c)]\right\}$
as the set of (mixed) actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no ambiguity, I abuse notation to (1) use $f(\alpha,\gamma)v(\alpha,\gamma)$ to mean the expected payoff given lotteries $\alpha\in\Delta(A)$, $\gamma\in\Delta(C)$, (2) identify actions $a,c$ with degenerate lotteries in $\Delta(A)$, $\Delta(C)$ respectively, and (3) use $a^{*}(c)$ to mean an arbitrary selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption the long-run player has a dominant stage-game action, labeled at $c=0$, and that they benefit from commitment.
@@
For all $a$, $\{0\}=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)$. Moreover,
$0\not\in\operatorname*{\arg\!\max}_{c\in C}\min_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
The first part of the assumption implies the set of payoffs attainable in the one-shot game can be represented as $\{f(\alpha,0)v(\alpha,0)\}_{\alpha\in\mathcal{N}(0)}$, i.e. the profiles where the long-run player chooses 0 and short-run players best reply. This simplifies the exposition by simplifying the “worst” punishments from deviating from a prescribed equilibrium and substantially simplifies the notational overhead. Towards this end, let $\overline{V}_{0}=\max_{\alpha\in\mathcal{N}(0)}\frac{f(\alpha,0)v(\alpha,0)}{1-f(\alpha,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}$ be their smallest such payoff.